Find the equivalent fraction of 3 / 5 having
numerator 27

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

We require numerator 27
Let N be the denominator of the fraction
∴27 / N = 3 / 5
3 × N = 5 × 27
N = (5 × 27) / 3
= 45
Therefore the required fraction is 27 / 45

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
The equivalent fraction of 3/5 having numerator 27 can be calculated as,
3/5 = 27/X
3X = 27×5
X = 135/3
X = 45
 
Hope it helps!
Thank you ✌️